OpenStudy (anonymous):

Solve the equation. |x| = 8 Choose the right answer please. a. {64} b. {-8, 8} c. {-8} d. {8}

OpenStudy (kira_yamato):

b

OpenStudy (jamesj):

Tell me, what is |-1|? or |-8|?

OpenStudy (alfie):

B. As a matter of fact, |-8| turns into 8 and 8 is 8 ;)

OpenStudy (anonymous):

c and d

OpenStudy (anonymous):

Are you sure shinigami1m? only i can choose one

OpenStudy (jamesj):

So let's go back to definitions. What is |-8| ?

OpenStudy (jamesj):

|-1| = 1 |-2| = 2 |-1/2| = 1/2 ... so |-8| = what?

OpenStudy (anonymous):

Question 4 Solve the equation. |x| = 8 Choose one answer. a. {64} b. {-8, 8} c. {-8} d. {8}

OpenStudy (jamesj):

Yes. So what is |-8| equal to?

OpenStudy (jamesj):

If you can't answer the question what is the value of |-8|, you can't answer your question, that's why I'm asking it.

OpenStudy (jamesj):

|-1| = 1 |-2| = 2 |-3| = 3 |-4| = 4 so |-8| = what?

OpenStudy (anonymous):

The question is the same. theri no other explanation.

OpenStudy (anonymous):

Only know that is not the (A) =64

OpenStudy (jamesj):

Ok horse. Last chance. Do you want to drink? What is the value of |-8| ?

OpenStudy (anonymous):

James the question does not have more information.

OpenStudy (jamesj):

|-1| = 1 |-2| = 2 |-3| = 3 |-4| = 4 |-5| = 5 |-6| = 6 |-7| = 7 |-8| = 8 Hence x = -8 is one value that satisfies the equation |x| = 8. Therefore either B or C is the answer. Now the question is, does x = 8 also work. I.e., is it the case that |8| = 8?

OpenStudy (anonymous):

I think it is 8.

OpenStudy (jamesj):

What is the definition of this | | sign? What does |x| mean?

OpenStudy (jamesj):

i know. I'm making sure you do.

OpenStudy (anonymous):

Not sure,but how can i do...

OpenStudy (jamesj):

|x| = x, if x > 0 |x| = 0, if x = 0 |x| = -x, if x < 0 Hence |1| = 1, because 1 > 0 |2| = 2, because 2 > 0 but |-1| = -(-1) = 1, because -1 < 0 |-2| = 2, because -2 < 0

OpenStudy (jamesj):

Similarly |-3| = 3 and |3| = 3

OpenStudy (jamesj):

So what is the value of |-4| ?